OpenStudy (anonymous):

Hello, Could somebody guide me to verify: sinx(tanx+1/tanx)=secx

OpenStudy (jdoe0001):

\(\bf tan(\theta)=\cfrac{sin(\theta)}{cos(\theta)} \\ \quad \\ \quad \\ sin(x)\left[tan(x)+\cfrac{1}{tan(x)}\right]\implies sin(x)\left[\cfrac{sin(x)}{cos(x)}+\cfrac{1}{\frac{sin(x)}{cos(x)}}\right] \\ \quad \\ sin(x)\left[\cfrac{sin(x)}{cos(x)}+\cfrac{cos(x)}{sin(x)}\right]\implies ?\)

OpenStudy (anonymous):

\[\frac{ \sin ^{2}(x)}{ \cos(x) } +\frac{ \cos(x)\sin(x) }{ \sin(x) } ?\]

OpenStudy (jdoe0001):

well.. is easier if you were to add the fractions inside first

OpenStudy (jdoe0001):

\(\bf sin(x)\left[tan(x)+\cfrac{1}{tan(x)}\right]\implies sin(x)\left[\cfrac{sin(x)}{cos(x)}+\cfrac{1}{\frac{sin(x)}{cos(x)}}\right] \\ \quad \\ sin(x)\left[\cfrac{sin(x)}{cos(x)}+\cfrac{cos(x)}{sin(x)}\right]\implies sin(x)\left[\cfrac{sin^2(x)+cos^2(x)}{cos(x)sin(x)}\right] \\ \quad \\ {\color{brown}{ recall\implies sin^2(\theta)+cos^2(\theta)=1}}\qquad thus \\ \quad \\ sin(x)\left[\cfrac{sin^2(x)+cos^2(x)}{cos(x)sin(x)}\right]\implies \cancel{ sin(x) }\left[\cfrac{1}{cos(x)\cancel{ sin(x) }}\right]\)

OpenStudy (anonymous):

thank you so much, you were really helpful

OpenStudy (jdoe0001):

yw